# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ 4.3 Goal-Centered Reward Design
Open-world targets often exhibit complex, unpredictable behaviors, requiring agents to adapt to varied movements and orientations. However, existing methods assume the drone is parallel to the target, leading to reward functions that inaccurately reflect tracking performance across different behaviors and perspectives (see Appendix C for theoretical proof). To address this, our reward function is defined only at the image level, relying on the target position within the image $I\in\mathbb{R}^{84\times 84}$. The reward decreases as the target moves away from the center $I_{O}$ of the image, and no reward is given if the target is outside as:
@@
$\small r_{t}=\left\{\begin{array}[]{lr}\tanh{(\alpha(1-x)^{3})}\ \ I_{target}\in I_{clip}\\ 0\ \ otherwise\end{array},\right.$ (2)
@@
where $x=\frac{\mid p_{vc}-I_{OG}\mid}{\mid E_{G_{c}}-I_{OG}\mid}$ represents the position of the target relative to the center of the image, which is calculated by projecting the image onto the ground. Specifically, we calculate the ratio of the distance from the projection $I_{OG}$ of the image center on the ground to the center $p_{vc}$ of the target (line segment $\mid\!p_{vc}\!-\!I_{OG}\!\!\mid$) and the distance from the intersection $E_{G_{c}}$ of the line connecting these two centers with the edge of the image projection to the center $I_{OG}$ (line segment $\mid\!E_{G_{c}}\!-\!I_{OG}\!\!\mid$), as shown in Fig. 4(b). The attenuation degree of $r_{t}$ can be adjusted using the hyperparameter $\alpha$, set to 4. The $\tanh(\cdot)$ provides a strong indication of the task target due to its relatively quick decay at the center of the image. $I_{clip}$ is the truncated image range set to prevent the drone from keeping the target at the edge of the image for a long time. The truncation of the image can be controlled using the hyperparameter $\lambda_{clip}$ as:
$\small\lambda_{clip}=\frac{W_{I_{clip}}}{W},$ (3)
where $W$ and $W_{I_{clip}}$ are the widths of the original image and the truncated image, respectively, we set $\lambda_{clip}=0.7$.
The reward function relies on the judgment of the corner points of the image: Left Up (LU), Left Down (LD), Right Up (RU), and Right Down (RD) and the target coordinates along with their projections. To map corner points onto the ground, the effective focal length of the camera ($f$) must first be estimated, as shown in Fig. 4(a). This can be calculated using the pinhole imaging principle [8], as follows:
$\small f=\frac{W}{2\tan{(\frac{1}{2}FoV)}},$ (4)
where $W$ is the camera image width (in pixels) and $FoV$ is the field of view of the tracker camera. After obtaining the focal length, we project the image corner points onto the target plane to check if the target falls within the image range, as shown in Fig. 4(a). This involves finding intersections between lines extending from the camera optical center through its corner points and the ground plane.
[…]
$\small\left\{\begin{array}[]{lr}l_{LUC}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{H}\\[4.5pt] l_{LDC}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{-H}\\[4.5pt] l_{RUC}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{H}\\[4.5pt] l_{RDC}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{-H}\\[4.5pt] \end{array}\right.,$ (5)
[…]
In coordinate transformations, a homogeneous transformation matrix (HTM) can be used to map a plane into another coordinate system. Suppose the analytical equation of the original plane $P_{0}$ is given by $A_{0}x+B_{0}y+C_{0}z+D_{0}\!=\!0$. The transformed plane $P_{1}$ has the equation $A_{1}x\!+B_{1}y\!+C_{1}z\!+D_{1}\!=\!0$. The homogeneous vectors for the two planes are $p_{0}\!\!=\![\hat{n_{0}},\!D_{0}]$ and $p_{1}\!\!=\![\hat{n_{1}},\!D_{1}]$, where $\hat{n_{0}}$ and $\hat{n_{1}}$ are the normal vectors. The HTM $T_{01}$ from plane $P_{0}$ to $P_{1}$ as:
$\small T_{01}=\begin{bmatrix}R_{01}&t_{01}\\ 0&1\\ \end{bmatrix},$ (6)
where $R_{01}$ and $t_{01}$ represent the rotation matrix and translation vector, respectively. The normal vector $\hat{n_{1}}$ of $P_{1}$ can be expressed as: $\hat{n_{1}}^{T}\!\!=\!\!R_{01}\hat{n_{0}}^{T}$.
Assume that the points $\mathbf{p_{0}}=[x_{0},y_{0},z_{0}]^{T}$ and $\mathbf{p_{1}}=[x_{1},y_{1},z_{1}]^{T}$ lie on the planes $P_{0}$ and $P_{1}$, respectively. These two points satisfy as follows:
$\hat{n_{0}}\mathbf{p_{0}}+D_{0}$ $=0,$ (7)
$\hat{n_{1}}\mathbf{p_{1}}+D_{1}$ $=0.$
Based on coordinate transformation principles, the relationship between the two points can be derived as $\mathbf{p_{1}}=R_{ct}\mathbf{p_{0}}+t_{ct}$. Therefore, by solving the simultaneous equations, the expression for constant term $D_{1}$ can be obtained:
$\small D_{1}=D_{0}-\hat{n_{1}}t_{01}.$ (8)
[…]
$\small A_{G_{c}}x+B_{G_{c}}y+C_{G_{c}}z+D_{G_{c}}=0.$ (9)
[…]
where $LU_{G}$, $LD_{G}$, $RU_{G}$ and $RD_{G}$ represent the mapping points of $LU$, $LD$, $RU$ and $RD$ on the ground plane. The expressions of $t_{LU}$, $t_{LD}$, $t_{RU}$, and $t_{RD}$ are as follows:
$\small\left\{\begin{array}[]{lr}t_{LU}=D_{G_{c}}(A_{G_{c}}f+\frac{1}{2}B_{G_{c}}W-\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{LD}=D_{G_{c}}(A_{G_{c}}f+\frac{1}{2}B_{G_{c}}W+\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{RU}=D_{G_{c}}(A_{G_{c}}f-\frac{1}{2}B_{G_{c}}W-\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{RD}=D_{G_{c}}(A_{G_{c}}f-\frac{1}{2}B_{G_{c}}W+\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] \end{array}\right..$ (11)
@@
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
@@
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{307}_{\pm 124}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{239}_{\pm 134}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{203}_{\pm 119}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{181}_{\pm 116}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$
@@
AOT $48_{\pm 5}$ $0.24_{\pm 0.02}$ $9_{\pm 1}$ $0.06_{\pm 0.00}$ $52_{\pm 11}$ $0.25_{\pm 0.03}$ $52_{\pm 6}$ $0.28_{\pm 0.03}$ $48_{\pm 5}$ $0.24_{\pm 0.02}$ $49_{\pm 7}$ $0.24_{\pm 0.02}$
D-VAT $49_{\pm 9}$ $0.26_{\pm 0.04}$ $48_{\pm 8}$ $0.27_{\pm 0.03}$ $50_{\pm 14}$ $0.24_{\pm 0.06}$ $8_{\pm 1}$ $0.05_{\pm 0.00}$ $51_{\pm 14}$ $0.25_{\pm 0.06}$ $13_{\pm 1}$ $0.07_{\pm 0.01}$
R-VAT $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $\textbf{230}_{\pm 115}$ $\textbf{0.67}_{\pm 0.27}$ $\textbf{156}_{\pm 119}$ $\textbf{0.55}_{\pm 0.31}$ $\textbf{201}_{\pm 121}$ $\textbf{0.64}_{\pm 0.30}$ $\textbf{163}_{\pm 115}$ $\textbf{0.51}_{\pm 0.29}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$
Table 2: Results of within and across scenes on the DAT benchmark.
[…]
Therefore, the coordinates of the car $p_{vc}$, along with the four intersection points ($LU_{G}$, $LD_{G}$, $RU_{G}$, and $RD_{G}$) and the image center $I_{OG}$ can be utilized to calculate the reward.
5 Experiments
Figure 5: Reward values during the training phase of our method and SOTA methods across six scenes.
5.1 Experimental Settings
[…]
Metrics. We use cumulative reward ($CR=\sum_{t=1}^{E_{l}}{r_{ct}}$) and tracking success rate ($TSR=\frac{1}{E_{ml}}\sum_{t=1}^{E_{l}}{r_{dt}}\times 100\%$) to evaluate the agent performance. The $CR$ primarily measures the tracker’s ability to keep the target centered in the image, where $r_{ct}$ is the dense reward at time step $t$ from Eq. 2, and $E_{l}$ is the length of the episode. The $TSR$ measures the tracker’s ability to maintain the target within the field of view, where $r_{dt}$ is the sparse reward (See Appendix C) at time step $t$ and $r_{dt}=1$ means the target within the field of view, $E_{ml}$ is the maximum episode length. Additionally, we position each agent relative to the target at $[0,\frac{\pi}{2},\pi,\frac{3\pi}{2}]~{}\text{rad}$ for 10 episodes each, totally 40 episodes. The mean and variance of these results are then calculated as evaluation metrics for the specific map. The model’s final cross-scene and cross-domain performance are both obtained by averaging results across different scenes.
Baselines. We reproduce the two SOTA VAT methods: AOT [33] and D-VAT [18]. See Appendix D for details.
@@ 5.1 Experimental Settings
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$

### v2

@@ 4.2 Goal-Centered Reward Design
For drone tracking a ground-based target, our reward is designed to characterize the target’s position in the image and guide the drone to keep it centered. Therefore, we first need to select an appropriate distance metric to quantify the proximity between the target and the center in the image plane.
Since the drone typically captures images from a top-down perspective, the image plane is not parallel to the ground. Due to the affine transformation, the projection of the image plane becomes a trapezoid (see Fig. 4(b)), and the physical distance between the drone and the target cannot directly correspond to their pixel distance. Existing methods e2e ; dvat compute the Euclidean distance between the drone and the target, which may not accurately reflect their spatial relationship in the image plane.
To address this issue, we employ a deviation metric $\phi(\cdot,\cdot)$ to measure the distance between the target and the image center projection, as illustrated in Fig. 4(b). Specifically, given a target point $P_{g}$ and the image center projection $C_{g}$, the deviation metric is computed by
$\small\phi(P_{g},C_{g})=\frac{\mid P_{g}-C_{g}\mid}{\mid E_{g}(P_{g},C_{g})-C_{g}\mid},$ (1)
where $\mid\!\!P_{g}-C_{g}\!\!\mid$ denotes the distance from $P_{g}$ to $C_{g}$ and $\mid\!\!E_{g}(P_{g},C_{g})\!-\!C_{g}\!\!\mid$ represents distance from point $E_{g}(P_{g},C_{g})$ to the center. $E_{g}(P_{g},C_{g})$ is the intersection of the line connecting $P_{g}$ and $C_{g}$ with the projected image boundary, as shown in Fig. 4(b).
The deviation $\phi(\cdot,\cdot)$ is designed to ensure that targets inside the image are closer to the center than those outside, with contours shown in Fig. 4(b).
Figure 3: Diagram of reward acquisition.
Figure 4: Reward design analysis diagram.
Principles for Reward Design. The objective of the VAT task is to keep the target at the image center. Thus, the targets closer to the image center projection should get higher reward values. For deviation metric $\phi(\cdot,\cdot)$, the design principle of the reward function $\mathcal{R}_{\phi}(\cdot)$ is defined as:
$\small\!\!\!\!\forall P_{1},P_{2}\in\mathcal{W},\text{ if }\phi_{1}<\phi_{2},\text{ then }\mathcal{R}_{\phi}(\phi_{1})>\mathcal{R}_{\phi}(\phi_{2}),$ (2)
where $\mathcal{W}$ denotes the valid region with non-zero reward, and $\phi_{1}$, $\phi_{2}$ represents the deviation from the target point to the image center projection.
Goal-Centered Reward Function. Our reward $r_{gc}(\cdot)$ decreases as the target moves away from the projected image center $C_{g}$, and is zero if outside, as shown as follows:
@@
$r_{gc}(P_{g}){=}\left\{\begin{array}[]{lr}\tanh{(\alpha(1{-}\phi(P_{g},C_{g}))^{3})},\ \ P_{g}\in\mathcal{I}_{clip}\\ 0,\ \ otherwise\end{array}.\right$. (3)
@@
The attenuation degree of $r_{gc}(\cdot)$ can be adjusted using the hyperparameter $\alpha$, set to 4. The $\tanh(\cdot)$ provides a strong indication of the task goal due to its relatively quick decay at the image center. $\mathcal{I}_{clip}$ is the truncated image range set to prevent the drone from keeping the target at the edge of the image. The truncation of the image can be controlled using the hyperparameter $\lambda_{clip}$ as: $\lambda_{clip}=H_{\mathcal{I}_{clip}}/H$, where $H$ and $H_{\mathcal{I}_{clip}}$ are the heights of the original and the truncated image. We set $\lambda_{clip}=0.7$.
Algorithm 1 Curriculum-Based Training (CBT)
1:Input: Initial policy parameters $\theta_{0}$, phase threshold $\eta$, total steps NN, rollout steps $n$
2:Initialize: Training phase $phase\leftarrow 1$, reward buffer $\mathcal{B}\leftarrow\emptyset$, rollout buffer $\mathcal{B}_{r}\leftarrow\emptyset$
3:for each step k=0,1,,N1k=0,1,...,N-1 do
4:  if $phase=1$ then
5:   Configure simple environment: linear target trajectories + no obstacles
6:  else
7:   Configure complex environment: varied target movements + obstacles/occlusions
8:  Collect transition $\tau_{k}=(s_{t},s_{t+1},a_{t},r_{t})$ with rewards calculated via (3)
9:  Append to buffer: $\mathcal{B}\leftarrow\mathcal{B}\cdot r_{k}$, $\mathcal{B}_{r}\leftarrow\mathcal{B}_{r}\cdot\tau_{k}$
10:  if $k\!\mod n=0$ then
11:   Update policy using PPO: $\theta_{k+1}\leftarrow\text{PPO\_Update}(\theta_{k},\mathcal{B}_{r})$
12:   Clear rollout buffer: $\mathcal{B}_{r}\leftarrow\emptyset$
13:  if $phase=1$ and $\frac{1}{|\mathcal{B}|}\sum_{r_{t}\in\mathcal{B}}r_{t}\geq\eta$ then
14:   Switch training phase: $phase\leftarrow 2$
15:   Clear buffer: $\mathcal{B}\leftarrow\emptyset$, $\mathcal{B}_{r}\leftarrow\emptyset$
More details about the Goal-Centered Reward. The reward function (Eq. 3) relies on the projections of the four corners and image center to compute deviation $\phi(\cdot,\cdot)$. As shown in Fig. 4(a), in the camera frame $\{\text{c}\}$, the image center and four corner points have the coordinates C(f,0,0)C(-f,0,0), $LU(\!-\!f\!,\!-\!\frac{1}{2}W\!,\!\frac{1}{2}H\!)$, $LD(\!-\!f\!,\!-\!\frac{1}{2}W\!,\!-\!\frac{1}{2}H\!)$, $RU(\!-\!f\!,\!\frac{1}{2}W\!,\!\frac{1}{2}H\!)$, $RD(\!-\!f\!,\!\frac{1}{2}W,\!-\!\frac{1}{2}H\!)$, where $W$ and $H$ are the image width and height and $f$ denotes the camera focal length, which can be computed using the pinhole imaging principle chartier2005introduction as: $f=\frac{W}{2\tan{(\frac{1}{2}FoV)}}$. $FoV$ is the camera field of view. Next, the equations of the lines connecting the image center and the four corner points to the optical center $O_{c}(0,0,0)$ can be obtained in frame $\{c\}$(light blue dashed lines in Fig. 4(a)):
$\small\left\{\begin{array}[]{lr}l_{LUO_{c}}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{H}\\[4.62497pt] l_{LDO_{c}}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{-H}\\[4.62497pt] l_{RUO_{c}}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{H}\\[4.62497pt] l_{RDO_{c}}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{-H}\\[4.62497pt] l_{CO_{c}}:y=0,z=0\end{array}\right$., (4)
where $l_{LUO_{c}}$ is the line connecting $LU$ to $O_{c}$, similarly for $l_{LDO_{c}}$, $l_{RUO_{c}}$, $l_{RDO_{c}}$ and $l_{CO_{c}}$. Thus, the projections of the points can be obtained by intersecting the lines with the ground plane.
Therefore, we next derive the expressions for the ground plane and the target. For clarity, we adopt a unified representation in frame $\{c\}$. In DAT scenes, the road surfaces are smooth. Thus, in the world frame $\{w\}$ (see Fig. 4(a)), the ground plane $G_{w}$ is defined as: $z=h$, where $h$ denotes the ground height. For simplicity, we here express $G_{w}$ in $\{c\}$ as $G_{c}\!:A_{g}x\!+\!B_{g}y\!+\!C_{g}z\!+\!D_{g}=0$, with $A_{g},B_{g},C_{g},D_{g}$ derived in Appendix C.2. Furthermore, the target coordinates $P_{v}=(x_{v},y_{v},z_{v},1)^{T}$ in $\{w\}$ can be transformed to $\{c\}$ using homogeneous transformation matrix HTM $T_{cw}$: $P_{g}=T_{cw}^{-1}P_{v}$.
Subsequently, the ground projections can be obtained by intersecting lines in Eq. 4 and $G_{c}$:
$\small\left\{\begin{array}[]{lr}LU_{g}:(-f,-\frac{1}{2}W,\frac{1}{2}H)t_{lu},\qquad&t_{lu}=D_{g}(A_{g}f+\frac{1}{2}B_{g}W-\frac{1}{2}C_{g}H)^{-1}\\[4.62497pt] LD_{g}:(-f,-\frac{1}{2}W,-\frac{1}{2}H)t_{ld},\qquad&t_{ld}=D_{g}(A_{g}f+\frac{1}{2}B_{g}W+\frac{1}{2}C_{g}H)^{-1}\\[4.62497pt] RU_{g}:(-f,\frac{1}{2}W,\frac{1}{2}H)t_{ru},\qquad&t_{ru}=D_{g}(A_{g}f-\frac{1}{2}B_{g}W-\frac{1}{2}C_{g}H)^{-1}\\[4.62497pt] RD_{g}:(-f,\frac{1}{2}W,-\frac{1}{2}H)t_{rd},\qquad&t_{rd}=D_{g}(A_{g}f-\frac{1}{2}B_{g}W+\frac{1}{2}C_{g}H)^{-1}\\[4.62497pt] C_{g}:(-\frac{D_{g}}{A_{g}},0,0)&\\[4.62497pt] \end{array}\right$., (5)
where $LU_{g}$, $LD_{g}$, $RU_{g}$, $RD_{g}$ and $C_{g}$ are the projections of LULU, LDLD, RURU, RDRD and $C$. Using target coordinates $P_{g}$ and Eq. 5, the reward is computed as Eq. 3. See Appendix C.2 for details.
4.3 Theoretical Gurantees on Reward Design
Existing methods e2e ; dvat assume a fixed forward camera view and use distance-based rewards. However, when the view changes, these rewards may fail due to the affine transformation effect in image projection. We hereby provide a theoretical analysis to show that commonly used distance-based rewards will fail when the camera deviates from a fixed horizontal forward view.
To this end, we define $\mathcal{R}_{d}(\cdot)$ as a distance-based reward using Euclidean distance between the target and the image center projection. A distance-based reward $\mathcal{R}_{d}(\cdot)$ satisfying Eq. 2 may still assign higher rewards to targets farther from the center under the metric $\phi(\cdot,\cdot)$, rendering it ineffective. In contrast, any deviation-based reward $R_{\phi}(\cdot)$ satisfying Eq. 2 can effectively reflect the target position.
Proposition 1
The commonly used Euclidean distance $d(\cdot,\cdot)$ between the target and the image center proposition does not align with the deviation $\phi(\cdot,\cdot)$ of the target from the image center projection, when the camera is not at a fixed horizontal forward viewpoint. That is:
$\small\!\!\!\!\exists P_{1},P_{2}\in\mathcal{I}_{p}\text{, s.t. }\phi_{1}\!<\!\phi_{2}\text{, }d(P_{1},\!C_{g})\!>\!d(P_{2},\!C_{g}),$ (6)
where $\phi_{i}=\phi(P_{i},C_{g})$, $P_{i}$ are points in the projection region $\mathcal{I}_{p}$, $C_{g}$ is the image center projection. See Appendix C.1 for theoretical proof.
Remark 1. A distance-based reward $\mathcal{R}_{d}(\cdot)$ satisfying Eq. 2 results in targets closer to the center receiving lower rewards, when the camera is not at a fixed horizontal forward viewpoint. That is:
$\small\!\!\!\!\exists P_{1},P_{2}\!\in\!\mathcal{I}_{p},\ s.t.\ \phi_{1}\!<\!\phi_{2},\ \mathcal{R}_{d}(d_{1})\!<\!\mathcal{R}_{d}(d_{2}),$ (7)
where $d_{i}=d(P_{i},C_{g})$, and $\phi_{i}=\phi(P_{i},C_{g})$ . This illustrates the failure of the distance-based reward under these viewpoints. See Appendix C.1 for theoretical proof.
Qualitative Analysis. According to the Theoretical Analysis above, rewards should decrease monotonically along the deviation contours in Fig. 4(b) as the target moves toward the projection boundary. Thus, the reward contours must align with the deviation contours. The contours of $r_{gc}(\cdot)$ in Fig. 4(b) perfectly align, indicating accurate position feedback. In contrast, D-VAT dvat (see Fig. 4(a)) shows misaligned contours, explaining its failure as noted in Remark 1.
4.4 Training with Curriculum Learning
DAT scenes contain numerous dynamic targets and obstacles, hindering convergence and performance. Progressively training the agent from simpler to more complex environments enhances performance and accelerates learning for the final task wang2021survey . Therefore, we propose a Curriculum-Based Training (CBT) strategy to optimize reinforcement learning training in complex environments.
To address the challenges, we employ the Proximal Policy Optimization (PPO) ppo algorithm, known for its efficiency in control tasks. To further enhance agent adaptability and robustness, we apply domain randomization during agent training. Specifically, we randomize the drone’s initial position and orientation relative to the target to promote diverse behaviors. Additionally, we randomize the gimbal pitch angle to improve the agent’s spatial perception. See Appendix C.2 for further details.
Given the scene complexity, we adopt a CBT strategy, which divides the model training into two stages. The first stage consists of a simplified environment with straight line target trajectories and no obstacles. The agent learns to center the target through the reward $r_{t}$ in Eq. 3. In the second stage, the agent encounters more varied target movements and complex visual information, such as tree occlusions and crosswalks. The goal of the agent is to develop stronger generalization abilities based on task understanding in the first stage. See Algorithm 1 for the pseudocode of the CBT strategy.
5 Experiments
5.1 Experimental Settings
Experimental Setup. We conduct cross-scene and cross-domain tests. The former tests an agent trained under daytime conditions in unseen scenes with the same weather. The latter evaluates the agent in the same scene under varying weather conditions. See Appendix E.1 for details.
Metrics. We use cumulative reward ($CR=\sum_{t=1}^{E_{l}}{r_{gc}}$) and tracking success rate ($TSR=\frac{1}{E_{ml}}\sum_{t=1}^{E_{l}}{r_{dt}}\times 100\%$) to evaluate the agent performance. $CR$ primarily reflects how well the agent centers the target over episode length $E_{l}$, while $TSR$ measures the ability to keep the target in view, with $r_{dt}=1$ meaning the target is within the view (See Appendix C), and $E_{ml}$ denoting the maximum episode length. Agents are initialized at four relative angles to the target ($0,\frac{\pi}{2},\pi,\frac{3\pi}{2}]$ rad), with 10 episodes per angle (40 total). The mean and variance of these results are calculated for each map, and the final cross-scene and cross-domain performance are averaged across different scenes.
Baselines. We reproduce two SOTA methods: AOT e2e and D-VAT dvat . Both baselines and other methods zhong2024empowering ; cvat use distance-based rewards. As concluded in Section 4.3, they may fail in tilted top-down views. Thus, these baselines sufficiently highlight GC-VAT superiority. See Appendix D.
Table 2: Results of within and across scenes on DAT benchmark.
@@ 5.1 Experimental Settings
CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR
@@
Ours $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{307}_{\pm 124}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{239}_{\pm 134}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{203}_{\pm 119}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{181}_{\pm 116}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$
@@
AOT $48_{\pm 5}$ $0.24_{\pm 0.02}$ $9_{\pm 0}$ $0.06_{\pm 0.00}$ $52_{\pm 11}$ $0.25_{\pm 0.03}$ $52_{\pm 6}$ $0.28_{\pm 0.03}$ $48_{\pm 5}$ $0.24_{\pm 0.02}$ $49_{\pm 7}$ $0.24_{\pm 0.03}$
D-VAT $49_{\pm 9}$ $0.26_{\pm 0.04}$ $48_{\pm 8}$ $0.27_{\pm 0.03}$ $50_{\pm 14}$ $0.25_{\pm 0.06}$ $8_{\pm 1}$ $0.05_{\pm 0.00}$ $51_{\pm 14}$ $0.25_{\pm 0.06}$ $14_{\pm 1}$ $0.07_{\pm 0.01}$
Ours $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $\textbf{229}_{\pm 115}$ $\textbf{0.67}_{\pm 0.27}$ $\textbf{156}_{\pm 119}$ $\textbf{0.55}_{\pm 0.31}$ $\textbf{201}_{\pm 121}$ $\textbf{0.64}_{\pm 0.30}$ $\textbf{163}_{\pm 115}$ $\textbf{0.51}_{\pm 0.29}$ $\textbf{162}_{\pm 106}$ $\textbf{0.54}_{\pm 0.26}$
Figure 5: Reward values during training.
Table 3: Results of cross domain on DAT.
@@
CRCR TSRTSR CRCR TSRTSR CRCR TSRTSR